Nathan made a triangular pennant for the band booster club. The are of the pennant is 80 square feet. The base of the pennant is 12 feet shorter than the height.
a) what are the lengths of the base and height of the pennant?
b)what are the dimensions of the pennant if the base is only 6 feet shorter that the height?
(I really do not understand this problem at all; do we use 1/2(b)(h)
please help soon!

Answer by checkley77(12844)

AREA OF A TRIANGLE=BH/2
(H-12)H/2=80
(H^2-12H)/2=80
H^2-12H=2*80
H^2-12H=160
H^2-12H-160=0
(H-20)(H+8)=0
H-20=0
H=20 IS THE HEIGHT.
20-12=8 FOR THE BASE.
PROOF:
8*20/2=80
160/2=80
80=80